A boat travels with the current at 23kph and travels against the current at 14kph. What is the rate of the still water? What is the rate of the current?
Answer by mananth(16949)

let still water rate of boat be x
and current rate be y
..
x+y=23
x-y=14
..
add the equations
2x=37
x= 18.5 speed of boat
y= 23-18.5=4.5 kph speed of current